NEPTUNE Canada

Announcements & Updates

New Website Launched

05-Sep-13

Our Oceannetworks.ca website has a new look, with the launch of our new brand and streamlined website. The new site includes content from our previous oceannetworks.ca, onccee.ca, VENUS and NEPTUNE websites. The former oceannetworks.ca and onccee.ca websites have now been taken offline, while the VENUS and NEPTUNE sites will remain online for a while longer. Eventually, all sites and data access facilities will be merged under oceannetworks.ca.

With the launch of this new website, we expect there to be some bugs and inconsistencies, which we're working to resolve. We welcome your feedback on the new site–let us know if you like what you see or if you have difficulties finding what you're looking for.

View the new website at: oceannetworks.ca


Exploration Now Live Event

21-Jun-13

Exploration Now, a new interactive project being led by Dr. Bob Ballard and his team at the University of Rhode Island, is providing live coverage of exporation from around the world, accessible to the general public at explorationnow.org.

Exploration Now has been streaming live video from the R/V Thompson for the past 2 weeks, while Ocean Networks Canada, a University of Victoria initiative, has been Wiring the Abyss in the NE Pacific.

Join a special broadcast by Dr. Robert Ballard today, 1:30 PDT, 4:30 EDT.


Wiring the Abyss 2013 Underway

30-Apr-13

Our Spring expedition season is now underway, aboard the CCGS John P. Tully. Our new Wiring the Abyss 2013 website provides access to photos, videos, stories and updates as the expedition progresses.

Follow the action via Twitter, Flickr, Facebook and Vimeo!


Folger Passage & Barkley Sound Workshop

23-Apr-13

On 28 May 2013, Ocean Networks Canada will host a workshop on research done at Folger Passage and Barkley Sound. It will bring together existing and new researchers to discuss present status, results and future opportunities for collaborative and cross-disciplinary projects.

More info


Live Video Now iPad Compatible

27-Feb-13

Today, we're happy to announce updates to our website enabling live video streams to iPad and iPhone devices. Please enjoy streaming video from our ReefCam and Barkley Canyon Axis, Upper Slope and Pod 4 cameras on your Apple mobile device! 


CFI Supports Expanded Research at Endeavour

15-Jan-13

Ocean Networks Canada is kicking off a year of exciting research and innovation at the Endeavour node of the NEPTUNE Canada subsea cabled network with a show of support from the Canadian Government.

Today’s $1.62 million funding commitment through the Canada Foundation for Innovation’s 2012 Leading Edge Fund is intended for future study and continuous monitoring of the Endeavour Segment.

Full story.


Mag 7.5 Earthquake Rocks Southeast Alaska

05-Jan-13

A magnitude 7.5 earthquake rocked the northeast Pacific Ocean just before 1 a.m. Saturday (PST), triggering a tsunami warning stretching from the southern coast of Alaska to the northern tip of Vancouver Island.

The quake centred about 150 kilometers northwest of Dixon Entrance, the maritime boundary between the U.S. and Canada.

More info


Job Opportunity: Administrative Coordinator

13-Dec-12

Ocean Networks Canada is looking for an administrative coordinator, responsible for day-to-day department-level administration of human resources-related activities.

Job Announcement


Earthquake Rumbles Recorded

02-Nov-12

Two NEPTUNE Canada hydrophones recorded the sounds of the recent magnitude 7.7 Haida Gwaii earthquake. The devices picked up the low-frequency rumbles from this earthquake, along with the calls of humpback whales in the area. The following recording is from a hydrophone on Barkley Upper Slope, positioned at a depth of 396 m. In this clip, sound has been sped up 250% to make the earthquake noise audible to the human ear. The squeaking noises prior to the main quake are the sped-up whale calls. (Headphones may be required to hear this recording clearly.)


Haida Gwaii Earthquake Sensed

29-Oct-12

A magnitude 7.7 earthquake struck southern portions of the Haida Gwaii archipelego the evening of 27 October 2012, generating a small tsunami. NEPTUNE Canada seismometers, bottom pressure recorders and hydrophones recorded the event, with the data providing new insights into near-shore tsunami propagation in coastal British Columbia.

Related Story


Live Video from the Reef!

23-Oct-12

We now have live video streaming from our Ocean Presence video camera on Folger Pinnacle. The camera is situated at a depth of 23 m and provides a stunning live view of the rich sea life living on the reef. Lights will be on all day today; starting tomorrow we'll have them on daily 8:00-8:45 AM and 12:00 - 12:45 PM Pacific Time.

ReefCam

Live Video Page for Scientists


Funding Secured!

03-Oct-12

Today, the governments of Canada and British Columbia announced a total of $41.7 million in new funding to support the operating costs of our parent organization, Ocean Networks Canada.

The funding is awarded through the Canada Foundation for Innovation’s (CFI) Major Science Initiatives Fund, which supports a portion of the operating costs of selected big science projects across Canada. Of the total, CFI is contributing $32.8 million and the BC government, $8.9 million.

Related links:


1 Year of Continuous Hydrothermal Vent Data

26-Sep-12

One year ago today the Cabled Observatory Vent Imaging System (COVIS) was powered back on after replacement of a malfunctioning extension cable connecting Endeavour Node to Main Endeavour Vent Field. (COVIS previously collected about a month of data before the cable failure and was then unpowered for 11 months while waiting for the cable to be replaced.) Now, COVIS has been collecting data continuously for 1 year since 26 September 2011 and has archived a once every 3-hour time series of acoustic imaging/Doppler data of the Grotto hydrothermal vent system. These data are now accessible on-line via Data Search.

rendering
Rendered image of Grotto hydrothermal vent and plumes generated from COVIS data, 26 September 2012.

The U.S. National Science Foundation has just recently funded a science team based at Rutgers University and the University of Washington Applied Physics Lab to begin a rigorous study of this data set (and future data) to further our understanding of the geophysical properties of hydrothermal vent systems and to further refine data processing techniques.


Ocean Webinar Series Announced

24-Sep-12

In a collaboration of the US National Science Foundation and the Group on Earth Observations, the NSF-funded Ocean Research Collaboration Network is proud to announce the first webinar in the series “Blue Marvel – Ocean Mysteries.” The series will look at the ocean and its impact on us - from life in the ocean to human life on Earth.

Registration required


3 Fall Workshops Slated

07-Sep-12

Three workshops, co-hosted by Ocean Networks Canada, are planned for November-December 2012. We invite your participation in:


Costa Rican Earthquake Detections

05-Sep-12

Today's magnitude 7.6 Costa Rican earthquake was detected by NEPTUNE Canada's ocean bottom seismometers. The earthquake occurred at 1442 UTC (7:42 PDT) and initial primary (or compressional) seismic P waves reached our network approximately 9 minutes later. These were followed, after 7 minutes, by S waves (secondary or shear waves). Finally, two varieties of surface waves were detected, Love waves at 1504 UTC (8:04 PDT) and Rayleigh waves at 1506 (08:06 PDT). Love waves are surface waves that cause horizontal shifting of the earth during an earthquake; Rayleigh waves are generated by the interaction of P and S waves at the surface.

All four wave arrivals show up clearly in the broadband seismometer data from ODP 1027 (second trace below), which is NEPTUNE Canada's deepest and quietest seismic location.

Today's earthquake is one of about 30 magnitude 6+ earthquakes to have struck the region over the past 40 years. A magnitude 6.5 tremor in 1973 claimed the lives of 26 people. USGS summary of this earthquake.

earthquake data
Detections of the magnitude 7.6 Costa Rica earthquake by NEPTUNE Canada instruments, 5 September 2012. (Click to enlarge.)
map
Locations of seismometers on the NEPTUNE Canada network. (Click to enlarge.)


Philippine Earthquake Detections

31-Aug-12

The magnitude 7.6 earthquake that the Philippines today also registered clearly in NEPTUNE Canada's ocean bottom seismometers. The earthquake occurred at 1247 UTC (5:47 PDT) and initial primary (or compressional) seismic P waves reached our network approximately 13 minutes later. These were followed, after 11 minutes, by S waves (secondary or shear waves). Finally the initial surface waves were detected at 1330 UTC (6:30 PDT). Surface waves take longer to arrive because they travel along Earth's surface rather than through Earth like P and S waves.

All three wave arrivals show up clearly in NEPTUNE Canada broadband seismometer data (top three traces below), but only the initial P wave arrival appears clearly in the scaled-up short-period seismometer data (4th trace below) because the S and surface waves are too long in period to be recorded by this device. Short-period seismometers are used to study small localized earthquakes, such as the thousands of small events that occur at Endeavour every year.

USGS summary of this earthquake

earthquake data
Detections of the magnitude 7.6 Philippines earthquake by NEPTUNE Canada instruments, 31 August 2012. (Click to enlarge.)
map
Locations of seismometers on the NEPTUNE Canada network. (Click to enlarge.)


Victoria Earthquake Detected

30-Aug-12

Although yesterday's magnitude 3 earthquake centred 20 km west-northwest of Victoria, BC occurred under land (26.5 km below the surface to be precise), several of our seafloor instruments detected the event. Small land earthquakes show up more clearly in land-based short-period seismometer data than they do in our broadband seismometers, which pick up a lot of ambient ocean noise.

Nevertheless, seismologists were able to tease out the earthquake signal. Plots from two seismometers are shown below; interestingly the more distant seismometer (ODP 1027) has a much clearer signal, due to much lower levels of ambient noise at that deep abyssal plain location. One of our seafloor hydrophones also recorded this event, as illustrated by the Folger Passage Hydrophone spectrogram inset on the map below.

earthquake data
Detections of the magnitude 3 Victoria earthquake by NEPTUNE Canada instruments, 29 August 2012. (Click to enlarge.)


Seeking a Director of Engineering

24-Aug-12

Competition is now open for a new Associate Director of Engineering at NEPTUNE Canada. Please let interested associates know about the job announcement.

Full job description


New Cruise Website

24-May-12

Wiring the Abyss 2012 is the name of our ocean expedition this year, set to depart on  27 May 2012. Follow our adventures via a new website we have created. Here, you can get all the latest video clips, photos, blog posts and ship updates. You can also follow the action underwater as ROPOS dives to the seafloor installing scientific instruments and observing fantastic sea creatures!


Marine Life Field Guide Released!

23-May-12

Our Marine Life Field Guide is now freely available as an iBook and a downloadable PDF. This guide is the first electronic and interactive visual reference to species of the deep sea. This unique new guide is a ‘living book’ that we'll be updating regularly with new creatures, images, videos and information.

Learn more.


Subsea Network Maintenance Today

09-May-12

Subsea network unavailable, beginning 1500 UTC, 9 May 2012 (0800 PDT). A full system outage is required to ensure the safety of repair personnel, as we work to diagnose power and communications problems at Folger Pinnacle instrument platform. The outage affects all underwater instruments; data collection will be stopped for the duration of the outage. This outage is expected to last until 0000 UTC, 10 May 2012.

Access to Oceans 2.0 and our data archives will not be affected.


News home.

DMAS & System Updates

Diving at Folger Pinnacle

17-Jul-13

16-17 July 2013 we are performing maintenance work at Folger Pinnacle, with the assistance of the Pelagic dive team.

16 July

Tasks completed:

  • Ocean Innovation Centre (ONCCEE) instrument platform recovered
  • CTD instrument recovered
  • New CTD instrument deployed
  • 3D camera cleaned

17 July

Planned tasks:

  • Inspection of the diver accessible portion of the cable running from node to Folger Pinnacle
  • Remediation of the cable wear points (if necessary) with split air hose
  • Redeployment of ReefCam
  • Installation of new marker buoy
  • Final cleaning of 3D camera lenses (time permitting)
  • Cleaning and application of anti-fouling (zinc) cream to ADCP transducers (time permitting)


Brief Software Outage this Afternoon

04-Jul-13

A routine maintenance update will be released for our Oceans 2.0 software this afternoon, beginning 13:30 PDT. No data interruption will be expected, however there will be momentary outage of the DMAS website as new code is deployed.


Update on Current Directions

27-Jun-13

Update 27 Jun 2013

The first phase of investigation has revealed that only VENUS data products are affected. We have confirmed the current directions delivered in NEPTUNE Canada Nortek Vector, HR AquaDopp and Vectrino data products are reliable.


CURRENT DIRECTIONS INCORRECT

26-Jun-13

Current Directions for Nortek Vector, HR AquaDopp and Vectrino are Incorrect

We have found a bug in one of the coordinate transformation programs used to deliver Nortek Vector, HR AquaDopp and Vectrino data. The current directions we are delivering in our data products are incorrect.

Current directions for all ONC (NEPTUNE and VENUS) Nortek Vectors, HR AquaDopps and Vectrinos are affected.

We are currently investigating the severity of this issue and we will post updates and notify our users with details when we have more information.

Thank you for your patience.


Archiving System Outage

21-Mar-13

Our "AD" archiving system will experience an outage 21 March 2013, 0900-1700 PDT, while a storage system "health check" is conducted. Archives will continue to be accessible via our University of Saskachewan mirror, however no new data will be ingested into the archives during the outage.


Unplanned Website Outage

22-Feb-13

Today our production websites and internal systems experienced an outage due to problems with the main disk storage system. The observatory wet plant, data collection and archiving were unaffected.

All systems have been restored and are now back online. We are consulting with vendors to address the causes of the outage.

We apologize for any inconvenience. 


Network Outage: 7 Feb 2013

04-Feb-13

Maintenance on our network switch handling research traffic for EDC2 will cause an outage on Thursday, 7 February 2013, 1230-1330 UTC (4:30-5:30 AM PST). The outage will affect access to NEPTUNE Canada, Ocean Networks Canada and ONC Centre for Enterprise and Engagement websites as well as VENUS data at the University of Victoria.

Data collection will not be affected at the shore stations, however database updates from non-NEPTUNE shore stations to our data archives will be delayed until after the outage is restored.


Server Outage: 16 Dec 2012

13-Dec-12

A website outage is planned for 16 Dec 2012, beginning 0700 PST (1500 UTC). The outage is expected to continue through most of 16 Dec 2012.

Our main database server will be upgraded to enhance website and data archive performance.

Affected systems:

Data collection will continue during the outage, with new data stored at our shore stations. These data will flow into the upgraded data archive server following resumed service.

For updates on this work, follow @neptune_systems on Twitter.


Repairs Complete, Everything Online

28-Nov-12

Repairs to the Folger Passage branchin unit are now complete and our subsea network is now completely back online. Please let us know of any problems you experience accessing data or devices on the network. 


Troubleshooting Continues

28-Nov-12

Our subsea network was powered up around midnight local time. We are troubleshooting communications with some of our infrastructure at Endeavour and Barkley Canyon.


Repairs Underway, Wet Plant Offline

26-Nov-12

Repairs continue on our subsea network. The Folger Passage branching unit is getting repaired to reenable bi-directional power/communications around the backbone. Wet plant outage to continue through today and possibly tomorrow as well.


Weekend Outage

23-Nov-12

Weather permitting, the entire NEPTUNE Canada wetplant will be shut down for repairs beginning Sunday November 25, 5:00 AM PST. Repairs to the Folger Passage branching unit will reenable bi-directional power and communications throughout the backbone cable. The outage is anticipated to last the entire day. Oceans 2.0 software and data access will be unaffected, although data collection will be suspended during the outage.

For updates, refer to the NEPTUNE Systems Twitter feed.


Node Outage

20-Oct-12

We are working to troubleshoot an outage at our Folger and ODP 1027 nodes. Communications with the science ports on those two nodes dropped and there have been some unusual power fluctuations. 


Software Update Today

10-Oct-12

A quick update to our Oceans 2.0 software will require a restart sometime this morning. Brief interruption expected.


Oceans 2.0 Software Upgrades

04-Oct-12

Our Oceans 2.0 data search and access software has been upgraded. New features include:

  • SeaTube has changed the way you can select high or low def videos. Up to three choices now, depending on the video.
  • Page added for sensor calibration. This is a sub-page of Device Details under Sensors.
  • Task Monitor shows details in the lower pane for tasks
  • New Workflow tab added to Device Details. Not being used yet for most instruments.

Release notes


DMAS Software Upgrading

05-Sep-12

Our Oceans 2.0 software is getting upgraded this afternoon - you may experience temporary slowness.


Software Updates

08-Aug-12

Our Oceans 2.0 software has been updated, including improvements to the metadata and access to seismometer health information. 

Release notes.


Software Updates Released

11-Jul-12

New updates on our Oceans 2.0 web software have been released, including a new vertical profile "cast" data product for instruments on POGO, the Profiling OceanoGraphic Observatory (formerly known as "VPS").

Release notes.


Potential Service Outage

10-Jul-12

We're upgrading our database storage array this morning. There is a small possibility of a short outage as we make the switch, and Oceans 2.0 users may be required to re-login.


New Software Release

24-May-12

Our Oceans 2.0 software has been upgraded. One new data product we look forward to seeing in action is the VPS Cast, which will display data through one full up-down deployment cycle of "Ogopogo", our vertical profiling system. Release notes are posted on our wiki.


Folger Communications Outage

18-May-12

A data outage for instruments connected to Folger Node is likely, beginning sometime today and extending into the middle of next week. The communications board for Folger Node has been exibiting intermittent problems and will be replaced next week, probably Wednesday 23 May or Thursday 24 May. Data archives are unaffected, but data collection from Folger Passage is likely to be interrupted.


Short Outage Today

17-May-12

A short communications outage on our is expected today sometime between 1100-1200 PDT, 17 May 2012. Data archives and web services will be unaffected.


Storage Upgrade Today

08-May-12
Today, we are upgrading storage for scientific data to 200TB. Archived data will continue to be available, but near-realtime updates to hydrophone and video data may be affected. Temporary outages are also possible.


Service Outage: 9 May 2012

07-May-12
Subsea network unavailable, beginning 1500 UTC, 9 May 2012 (0800 PDT) A full system outage is required to ensure the safety of repair personnel, as we work to diagnose power and communications problems at Folger Pinnacle instrument platform. The outage will affect all underwater instruments; data collection will be stopped for the duration of the outage. This outage is expected to last until 0000 UTC, 10 May 2012. Access to Oceans 2.0 and our data archives will not be affected.


Server Outage

03-May-12

Our Oceans 2.0 data search and analysis software suite is getting upgraded today and will be taken offline for about an hour beginning 13:00 UTC, Pacific Daylight Time. Data acquisition will not be affected during the outage.


Folger Pinnacle inactive until the summer repair season

27-Apr-12

The Folger pinnacle platform has been shut off due to an isolation fault. It will remain offline until after recovery and redeployment with an upgraded 3d camera in mid-summer.

On 2 April 2012 a persistent isolation leak of ~10mA was detected on 400V line to the Folger Pinnacle by the node. The junction box at Folger Pinnacle was not affected, which implies that the leak is occurring either along the cable or at the connectors between Folger Node and Folger Pinnacle Junction Box.

After consideration, the damage caused by a 4W leak was deemed too unpredictable to continue operating the instruments on the platform.

Our engineering team is working to procure new cables and connectors to be installed during our camera maintenance and redeployment operations this summer.


News home.

Cruise Notices

October 25th 21:00 PDT Update

25-Oct-13

On the last day of the Vector expedition, we have successuly deployed the repaired hydrohone array at 170 m in the Strait of Georgia, recovered autonomous mooring from the Saanich Inlet sill location, and recovered the remaining pig bones from Saanich Inlet.

Read more in Daily Log - October 25, 2013


October 24th 21:00 PDT Update

24-Oct-13

Today we successfully deployed the new forensic experiment, recovered both Benthic Boundary Layer (BBL) platform and the IOS hydrophone array. All operations were in the Strait of Georgia. Also, today it was decided not to redeploy the new hydrophone systems (ONC Innovations) that were recovered two days earlier. The hydrophones will require more time to fine-tune and will be re-deployed in 2014.

Read more in Daily Log - October 24, 2013


October 23rd 21:00 PDT Update

23-Oct-13

Day 3 of the expedition featured servicing and re-deployment of the Delta Dynamics Laboratory (DDL) platform, recovery of the Seismic Liquefaction In-situ Penetrometer (SLIP), and an attempt deploy the next forensic experiment in the Strait of Georgia, which was then postponed to the following day.

Read more in Daily Log - October 23, 2013


October 22nd 21:00 PDT Update

22-Oct-13

Another busy day aboard CCGS Vector has come and gone. The ONC engineering crew succesfully serviced and redeployed the core instruments platform in Saanich Inlet, retrieved forensic experiment platform and brought to shore two ONC Inovations' hydrophone arrays for the on-shore servicing.

Read more in Daily Log - October 22, 2013


October 21st 21:00 PDT Update

21-Oct-13

The first day of the VENUS observatory maintenance cruise aboard the Canadian Coast Guard Ship Vector is now complete. A week-long expedition began with servicing instruments at the Fraser River delta location of the the Strait of Georgia array of VENUS. 

Read more in Daily Log - October 21, 2013


CCGS Vector Expedition Begins

20-Oct-13

Between October 20-27, 2013 Ocean Networks Canada is conducting the last scheduled maintenance expedition of the 2013 calendar year. This cruise's focus is maintenance of the VENUS observatory installations from the Canadian Coast Guard Ship Vector using the CanPac Divers' ROV Oceanic Explorer. The expedition's objectives include:

  • Maintain and clean the core sensors on all VENUS instrument platforms (VIP)
  • Recover, service, and redeploy the Delta Dynamics Laboratory near Sand Heads
  • Recover the Seismic Liquefaction In situ  Penetrometer (SLIP) at the Delta site
  • Recover, service, and redeploy three hydrophone arrays in the Strait of Georgia
  • Recover (from Saanich Inlet, with bones), service, and redeploy (Strait of Georgia) the Forensics Platform, with two new pig carcasses.
  • Recover the Bottom Boundary Layer (BBL) platform from the Central Strait of Georgia site

Check the daily logs and cruise status sections for updates.


September 19th 1630 PDT Update

19-Sep-13

A very successful cruise aboard the R/V Falkor has come to an end. The ship docked in Victoria at Ogden Point around 0830 yesterday morning, and we spent the day off-loading equipment and cleaning the labs. The final day of operations was spent at Tully Canyon where we did two extensive survey dives to examine benthic life in the low-oxygen environment. In the evening, we had a barbecue all together on the upper deck of the ship to celebrate the end of the cruise. Many thanks to all those who contributed to make this expedition possible: Schmidt Ocean Institute, the crew of the R/V Falkor, the ROPOS team, our visiting scientists, and the staff supporting us on shore.


September 16th 1900 PDT Update

16-Sep-13

Operations in Barkley Canyon finished at noon today. The last dive, the 600 m transect, completed successfully after a couple of minor delays overnight. We found quite a bit of kelp which was sampled, in addition to sponges and anemones. Now we are doing a multibeam survey and looking forward to two dives starting at midnight at Tully Canyon, a short steam away across the U.S. border.

Read more at Daily Log - September 16, 2013


September 15th 1000 PDT Update

15-Sep-13

Dive R1655, the 300 and 400 metre deep transects of Barkley Canyon, took place overnight. It was very interesting - especially so for those of us who have seen the Ocean Networks Canada sites in Barkley Canyon many times - because of the rugged topography and prevalence of rocks and pebbles that we don't see further down in the canyon. Kelp found, samples taken, and we are now underway to the 1500 m transect site.

Read more in Daily Log - September 15, 2013


September 14th 1520 Update

14-Sep-13

The tether retermination went very quickly, however before we could get back in the water, a power issue with the telemetry can also needed to be resolved. ROPOS worked through the night, allowing us to start Dive R1564 early this morning. In that dive we completed a 200 m deep transect of the head of Barkley Canyon. We encountered a region where fragile pink sea urchins were incredibly abundant. We did suction samples of the sediment surface, push cores and gathered sea urchins and sea cucumbers for analysis. Dive R1655: 300 m and 400 m transects will be starting in a few hours.

Read more in Daily Log - September 14, 2013


September 13th 0945 Update

13-Sep-13

The first transect at 2000 m resulted in some interesting samples including sea pens, sea cucumbers and an anenome for which we currently have no identification. Unfortunately, last night on recovery there was a kink in the ROV tether which has forced us to pause survery operations to reterminate the umbilical. We hope to be diving again early this afternoon.

Read more in Daily Log - September 13, 2013


September 12th 1345 Update

12-Sep-13

This morning we finished the last instrument/experiment deployment dive of the cruise. Next on the programme are visual and sampling transects at staggered depths across Barkley Canyon to study the relationship between oxygen levels and observable fauna. We are at bottom, just about to start the first transect at 2000 m, below the oxygen minimum zone (Dive 1653).

Read more in Daily Log - September 12, 2013


September 11th 0830 PDT Update

11-Sep-13

Dive 1651 at Barkley Canyon POD 3 has just started. In this dive, we will be deploying the first algae treatment of a sediment enrichment experiment. We will be distributing phytodetritus within the frames that we placed in view of the camera yesterday. While waiting for the treatment to settle, we will survey at Wallyland and return to the Coral Cliff to recover the second ADCP.

Read more in Daily Log - September 11, 2013


September 10th 0530 PDT Update

09-Sep-13

The operations at Clayoquot Slope (formerly OPD 889) went well: we swapped the CTD for a CTD-O2, we exchaned SCIMPI's data logger and recovered the broadband seismometer auxiliary platform. Now, the Falkor has arrived at Barkley Canyon, and we are getting ready for a morning dive to install a sonar at Wallyland, clean the crawler and deploy the first stage of a sediment enrichment experiment at POD 3.  The dive will be starting around 0800.

Read more in Daily Log - September 9, 2013 and Daily Log - September 10, 2013


September 8th 1230 PDT Update

08-Sep-13

We have completed all our work in Saanich Inlet and the Strait of Georgia: push cores have been taken, 2 cliff climbs and a transect were completed, the Buoy Profiling System is connected and organisms have been sampled. We are now heading offshore to Clayoquot Slope where we will first perform a navigation system calibration and followed by instrument operations starting early Monday morning.

Read more in Daily Log - September 8, 2013


September 7th 0720 PDT Update

07-Sep-13

The first dive of Leg 2 of the Falkor cruise has started! We are in Saanich Inlet, starting a visual transect of Patricia Bay. The purpose of this dive is to discover the distribution of flatfish in relation to depth and oxygen levels. We will be collecting live samples of these fish for further analysis in the lab.

Read more in Daily Log - September 7, 2013


R/V Falkor Expedition, Leg 2 Begins

06-Sep-13

Today the R/V Falkor set sail from Ogden Point, kicking-off Leg 2 of the "Open Ocean to Inner Sea expedition". Preliminary tasks include several benthic surveys in Saanich Inlet, along with the physical connection of the new Buoy Profiler System recently installed in Saanich Inlet to the VENUS node in Patricia Bay. 

Read more in Daily Log - September 6, 2013

Read more about the expedition in "ONC and Schmidt Team up for Leading Edge Science Research".


R/V Falkor Expediton, Leg 1 is Complete

31-Aug-13

August 31, 15:00PDT Update

Today the R/V Falkor pulled into the Esquimalt Harbour and the first leg of the ONC-SOI 2013 joint expedition come to a conclusion. We packed up our personal possessions, cleaned the laboratories, and dismantled the scientific gear. By mid-afternoon the aft deck was clear and a truck-load of equipment was return to UVic’s Marine Technology Centre. Tomorrow, the loading of ROPOS begins.


August 30th 2300 Update

30-Aug-13

Prior to our most recent exposure to this special oceanographic system, MVP is most often a reference to Most Valuable Player. It is an award, often handed out annually in a sports league, to a player who has contributed the most in a team’s effort to achieve their goals. The MVP on this cruise was, in both senses, the Moving Vessel Profiler.

Read more in Daily Log - Aug 30, 2013


August 29th 2300PDT Update

29-Aug-13

As the penultimate science day, we are really ticking off the boxes of what needs to be accomplished. Yesterday we did a little run along the coastal front next to Vancouver Island. We then zig-zagged yet again along the shelf break, and finally we did a nice tidal time series over the Zeppelin Bank and across the Tully canyon over a flood tide. The final MVP survey tomorrow will be relatively short (12 hours), before we bang off our last CTD and head for home.

Read more in Daily Log - Aug 29, 2013


August 28th 2300PDT Update

28-Aug-13

Today the edits to the plan were minor, and we have started our final push. The weather is grey, with low clouds, irregular seas, and cool temperatures. The forecast calls for an increase in the SE wind. The MVP remains in the water nearly all day, with periodic checks on the fish to inspect for kelp and wear. Our adaptive sampling has a few days left to map some of the key features one last time.

Read more in Daily Log - Aug 28, 2013


August 27th 2300PDT Update

27-Aug-13

The bumpy sea abated through the night, and by morning, the winds were light and the sea a confused wash of old and dying waves. We steam along at our preferred MVP speed of eight knots, and sample with particular purpose. The MVP is used to find exact locations across the continental shelf front for our final set of CTD casts. We are very pleased with these samples. By the end of the day, we have started a marathon of an MVP survey, very likely our last hurrah. If we never give up, luck will find us.

Read more in Daily Log - Aug 27, 2013


August 26th 2300PDT Update

26-Aug-13

There has been a standing forecast the last couple of days for gale force winds. They started to build last night, and outside of Barkley Sound, we are seeing a steady 30 knots. Both the MVP and CTD seem game at first, so the sampling program should not be affected, but we will take a little more care with deck operations.

Read more in Daily Log - Aug 26, 2013


August 25th 2300PDT Update

25-Aug-13

Today a plan for the remainder of the cruise was prepared. We'll see how long it lasts, events change regularly at sea. But it’s a workable plan and includes the remaining cruise objectives. The adaptive science is relentless and gives me no quarter.

Read more in Daily Log - Aug 25, 2013


August 24th 2300PDT Update

24-Aug-13

The day has been a gentle, long sequence of repeated MVP profiles. Counting up into the hundreds, the ship steams along at eight knots, while the profiler surfaces and sinks, again and again.

Read more in Daily Log - Aug 24, 2013


August 23rd 2300PDT Update

22-Aug-13

We complete the second MVP survey along the centre of Tully Canyon, ending near the NEPTUNE Folger site and Bamfield. We let Stephane Gauthier off at Bamfield, so he can get ready for another cruise. We conduct a multi-beam survey of Folger pinnacle, and start the third MVP survey. The oxygen story is both coming into focus and confirming some recent findings by Bill Crawford about the presence of very low oxygen water in the central shelf west of Vancouver Island. At the end of the day, rather dazed and less confused, we see bubbles.

Read more in Daily Log - Aug 23, 2013


August 22nd 0900 PDT Update

22-Aug-13

Today we completed the second MVP survey along the Tully Canyon thalweg, ending near the NEPTUNE Folger site and Bamfield. We let Stephane Gauthier off at Bamfield.  We conducted a multi-beam survey of Folger pinnacle, and the third MVP survey.

Read more in Daily Log - Aug 22, 2013


August 21st 0900 PDT Update

21-Aug-13

Late in the evening Tuesday we started the first moving vessel profiler (MVP) survey. It was completed this morning, and has given us some food for thought. The middle of the day was devoted to calibrating the ship’s new ultra-short baseline (USBL) transponder, before we started an eagerly anticipated MVP transect up the Tully Canyon.

Read more in Daily Log - Aug 21, 2013


August 20th 0900 PDT Update

20-Aug-13

The day has been a very typical oceanographic sampling day, dominated by “classic hydrography”, which is the measurement of water properties as a function of depth at fixed stations across the ocean. Continuing from yesterday, we completed the initial CTD and Rosette survey, including portions of the Barkley Canyon and La Perouse C lines.

Read more in Daily Log - Aug 20, 2013


August 19th 0800 PDT Update

19-Aug-13

This morning at 0800 we arrived at the entrance to Juan de Fuca Strait. A light fog, but relatively calm seas. The morning briefing on the bridge confirms today’s simple schedule: a final configuration of the MVP and then we start an intense CTD/Rosette survey.

Read more in Daily Log - Aug 19 2013


August 18th 09:00 PDT Update

18-Aug-13

At 09:00 PDT R/V Falkor slipped the dock and headed north out of Nanaimo into the central Strait of Georgia. First step is to work out sampling procedures and calibrate ecosounder.

Read more in Daily Log - Aug 18 2013


August 17th 16:00 PDT Update

17-Aug-13

Still at dock in Nanaimo. It has been a busy day aboard the R/V Falkor. We have been both physically and virtually preparing for our departure at 0900 August 18.

Read more in Daily Log - Aug 17 2013


R/V Falkor Cruise Begins

16-Aug-13

Our fourth major expedition of Summer 2013 will be underway on August 18, embarking from Nanaimo, BC on board the Schmidt Ocean Institute's R/V Falkor. We invite everyone to share in daily expedition operations and adventures, via Wiring the Abyss 2013 and Twitter - #Abyss13.

Read more in Daily Log - Aug 16 2013


R/V Thompson Expedition Complete

27-Jun-13

We have now completed our expedition aboard the R/V Thompson and demobilized from ship. We're looking forward to our next expedition about the R/V Falkor, beginning August 17.


June 25th 00:25 PDT Update

24-Jun-13

We are homeward bound! The R/V Thompson is presently in the Strait of Juan de Fuca, on its way to port in Seattle. Our scheduled ETA is 1100, June 25th, following which trucks will arrive to load our gear, and the ONC staff will return to Victoria via the Clipper. The last stop at Clayoquot Slope (formerly ODP 889) was short, but we nevertheless did 2 dives. We recovered the CSEM instrument, investigated a vent hole and visited the SHRIMPI and SCIMPI instruments.


June 23rd 23:45 PDT Update

23-Jun-13

Dive M0017 has come and gone. We deployed four instruments at Cascadia Basin (formerly ODP 1027), our deepest site at 2660 m. We are now steaming to Clayoquot Slope (formerly ODP 889) to recover the CSEM instrument, investigate a vent hole, and collect data from SCIMPI and Shrimpi instruments.  This will be our last stop of the cruise before returning to port in Seattle.


June 23rd 09:00 PDT Update

23-Jun-13

Endeavour is behind us and we have moved on to Cascadia Basin (formerly ODP 1027). At this site, we will deploy a low-frequency (LF) hydrophone, a bottom pressure recorder (BPR) experiment, an acoustic Doppler current profiler (ADCP) and a conductivity-temperature-depth (CTD) instrument with an oxygen sensor. Dive M0017 has just started. The ROV and the instrument package on the ship's wire are descending and will be on bottom shortly.


June 22nd 17:30 Update

22-Jun-13

The cable spool was recovered this morning and is now on deck. Dive M0016 at Main Endeavour Field is underway. The deployment of the Remote Access water Sampler (RAS) just came to a successful end. Next, we are taking some tubeworm samples, and then recovering the ROV. Next stop, Cascadia Basin (formerly, ODP 1027).


June 22nd 01:00 PDT Update

21-Jun-13

We have moved back to the Endeavour Mothra site to recover the Mothra cable and LOCO cable laying system. Dive M0015 has just begun. This afternoon, we dove at the Endeavour North location (RCM-North) to continue the diagnosis of the communications failure. We determined that plugging the existing cable into a different node port did not bring the site online. After performing several other tests, it is suspected that the cable might be at fault for the problem.


June 21st 07:20 PDT Update

21-Jun-13

At the end of Dive M0011, the ROV tether caught on some rocks around the base of Grotto which necessitated maintenance after recovery to the spooling system. Dives M0012 and M0013 also had to end early due to related problems. The deployment of the Remote Access water Sampler (RAS) was interrupted.  We are taking a break from Main Endeavour Field to continue the diagnosis of the communications issues at the RCM-North site, but will return later to finish MEF operations.


June 20th 07:50 PDT Update

20-Jun-13

Dive M0011 is coming to an end. We need to recover the ROV in order to reset the gas tights for RAS (remote access water sampler) deployment. In the last day, we completed Tempo-Mini's deployment, a visual survey of Grotto for photogrammetry, deployment of temperature loggers near Tempo-Mini and we designed and lowered new bio-buckets for tubeworm collection, which will occur on the next dive. RAS will be lowered on the ship's wire this morning.


June 19th 16:00 PDT Update

19-Jun-13

Dive M0011 has now been underwater for nearly 40 hours. The crew, ROV pilots, science and engineering teams are working hard to keep up with 24 hour operations. Tempo-Mini deployment is ongoing; the ROV is getting ready to position the instrument on Grotto. Today is a beautiful sunny day and we are still fortunate with calm seas.


June 19th 01:10 PDT Update

18-Jun-13

Dive M0011 ongoing and proceeding well. Following the last update, BARS has been successfully deployed and water samples and temperature profiles for COVIS were completed. We also took tubeworm samples from North Tower and gas tights for BARS calibration. At the time of this writing, Tempo-Mini is waiting for deployment on the ship's wire and will be placed next. Other upcoming tasks: temperature logger replacement, photogrammetry survey, short-period seismometer and auxilliary platform deployment, and further sampling.


June 18th 09:20 PDT Update

18-Jun-13

Dive M0011 at Main Endeavour Field in progress. Instruments are being lowered to the seafloor on a specially built platform on the ship's wire and guided to the bottom by the ROV. The first load included the 600 kHz ADCP (Acoustic Doppler Current Profiler) and BARS (Benthic and Resistivisty Sensors) vent probe. Early this morning, the ADCP was deployed. We are currently in the process of placing BARS in a vent on the Grotto crown.


June 17th 19:00 PDT Update

17-Jun-13

We are moving on station to begin dive operations at Main Endeavour Field. In the next two days, we will deploy RAS, BARS, Tempo-Mini, a 600kHz ADCP, temperature loggers, and do biological and visual sampling. We anticipate being on bottom around midnight tonight, PDT, but it will still take us some time to begin operations, so the first deployments will start in the early morning.  We left RCM-North without successfully diagnosing the failure. The spare junction box did not resolve the issue; we will return later in the cruise.

 


June 17th 00:00 PDT Update

16-Jun-13

We have moved on to the RCM-North site to diagnose the communications failure at that location. Dive M0010 is in progress. The first test is to evaluate the functioning of the junction box. We have lowered a spare instrument platform on the ship's wire and we will be attempting to connect it shortly. The attempts to mate the connector to LOCO on previous dives were not successful and therefore we were unable to lay the Mothra-Midpoint cable up to now.


June 16th 14:00 PDT Update

16-Jun-13

Dive M0009 in progress. We attempted to bring LOCO, the cable laying system, up to 250 m in order to try to mate the connector to the cage. We were able to plug it in, but the connector would not stay mated. The ROV is presently on the way back to the bottom with the spool. We will not be able to lay the cable at this time, however, we are in contact with Oceaneering and they are working on a solution. In the meantime, if we can't lay the cable, we intend to go to RC-North to diagnose the failure at that site.


June 16th 00:30 PDT Update

15-Jun-13

Dive M0008 is underway. The goal of this dive is to bring the cable layer (LOCO) attached to the Mothra cable spool up to a shallower depth at which we hope it will be possible to mate the connector between the cage and LOCO. Our tests on Dive M0007 revealed that the dummy connectors could be successfully mated between 250 and 500 m depth. When connected, we will return the spool to the seafloor and begin the Mothra-Midpoint cable lay. Once started (at the earliest, Sunday morning PDT), the lay will take approximately 12 hours to complete.


June 15th 09:40 PDT Update

15-Jun-13

More dive woes overnight. We had to do an emergency recovery because the pan & tilt was damaged on the main HD camera during dive M0006 before the connection to LOCO could be established. Weather is still co-operating. All hoping for good news later today.


June 14th 20:30 PDT Update

14-Jun-13

The ROV is back on deck again. We were unable to start the cable lay because of problems mating the connector to power the cable laying system. The ROV team is working on the issue, and we should be diving again soon.


June 14th 11:30 PDT Update

14-Jun-13

We are beginning the Mothra-Midpoint cable lay. The heading required for the cable lay might interfere with satellite communications from the ship. (There is a small "block zone" for visibility of the satellite with the antenna on the ship.)  If there are outages in the live feed, don't panic and please check back later!


June 14th 09:30 PDT Update

14-Jun-13

Dive M0005 is now underway. The end of the cable has been connected to the Mothra Instrument Platform; cable lay to Midpoint upcoming. The last operation of Dive M0004 was to recover four acoustic releases which will sit between the cage and the spool to provide a release mechanism should the cable lay be interrupted. M0004 completed around 04:30 this morning.


June 14th 00:00 PDT Update

13-Jun-13

Dive M0004 is coming to an end. The spool of cable for the Mothra-Midpoint cable lay has been deployed on the ship's wire and readied by the ROV for the cable lay. We will perform the lay on the next dive overnight and into the morning. Weather is perfect for operations: flat seas.


June 13 09:25 PDT Update

13-Jun-13

Dive M0003 has just started; dive objective to lay Mothra to Midpoint cable. M0002 was another short dive, following which several issues were resolved overnight with the ROV and navigation. The weather and seastate are perfect for cable laying so we are optimistic about upcoming operations.


June 13th 00:15 PDT update

12-Jun-13

We are currently at Mothra running a CTD. A first test dive (M0001) with the ROV revealed some minor problems which got sorted out and we are about to conduct another test before we go on the first science dive (Mothra).


June 12th 11:30 PDT update

12-Jun-13

We are approximately 2 hours away from Endeavour, heading to Mothra for our first operations.  We will start with a CTD cast at the site, and then evaluate weather conditions for starting the Mothra to Midpoint cable lay. Winds have been 15-20 kn this morning with a approximately 2 m seas.


June 11th 15:40PDT Update

11-Jun-13

We are currently underway, steaming through Puget Sound en route to the Strait of Juan de Fuca. From there, we will transit to the Endeavour site where we will start our maintenance operations. The ROV has completed two important tests: a successful wet test at the dock in Seattle overnight, and a navigation test this afternoon. We are taking advantage of the 28 hour transit to finish organizing ourselves onboard for the upcoming operations.


June 10th 10:00PDT Update

10-Jun-13
Mobilization is in progress aboard the R/V Thompson for the Ocean Networks Canada June Cruise. We anticipate being ready to sail Tuesday morning. The Oceaneering Millenium ROV is onboard and nearly ready for wet test. All hands are busy with setup of video transmission, network systems, instrument preparation, and securing the deck for transit. The rest of the science party will be arriving later this afternoon on the Clipper from Victoria. Stay tuned!


June 10th 08:00PDT Update

10-Jun-13
Welcome to the R/V Thompson cruise scheduled for June 11 to 26th, 2013. We are currently mobilizing in Seattle, Washington, with crew and visiting researchers arriving from around the world. This expedition will feature the Oceaneering ROV “Millenium,” operated by Canpac Divers. Check out our live video to see the "Millenium," and stay tuned for updates in this space!


May 27th to June 9th

29-May-13

No Expedition Operations at this time. 
Watch for our Thompson expedition - June 9 to June 26.


May 27th 10:00PDT Update

27-May-13

Our expedition aboard the CCGS Tully has concluded and we are now demobilizing from ship at the Institute of Ocean Sciences dock in North Saanich, BC.


May 25th 08:00PDT Update

25-May-13
We are currently transiting from the Strait of Georgia to Saanich Inlet. We will be diving this morning at 1000 to complete a series of cliff surveys in Squally Reach. This afternoon we will shift to our Saanich Node site for one final dive before coming alongside.


May 24th 09:00PDT Update

24-May-13

This morning we are sitting off the mouth of the Fraser River at our Delta Dynamics Laboratory site. We will be diving this morning to work on a hydrophone array. The outflow from the Fraser plus very large tides give us a very narrow time window in which to complete the dive. This afternoon we will shift to our East Node site for another dive.


May 23rd 07:00PDT Update

23-May-13

This morning we are in Saanich Inlet to recover the Micro Squid platform and replace a damaged oxygen sensor. We will work on the package and then redeploy it shortly before lunch. The Tully will then transit to the Strait of Georgia where we will dive to conduct maintenance on our hydrophone arrays.


May 22nd 16:00PDT Update

22-May-13

We have just completed a successful dive in Tully Canyon. We are now conducting a short sonar survey and will then begin the transit back to Saanich Inlet. This marks the end of the NEPTUNE maintenance operations for Tully Leg #2.

We will be diving at 0730 tomorrow morning in Saanich Inlet to conduct maintenance on the VENUS observatory.


May 22nd 07:00PDT Update

22-May-13

Winds have dropped but the seas are still too high in Barkley Canyon to dive.  We are heading to an alternate site in Tully Canyon where conditions look more favourable.

We should be diving at 0830.  Our plans are to spend the day at Tully Canyon and depart for Saanich Inlet in the late afternoon. 

Tomorrow we will  conduct maintenance dives on the VENUS observatory.  

 


May 21st 08:00PDT Update

21-May-13

Winds have been building steadily overnight and are now blowing from the north at 35 kts. Seas have increased and we have been forced to cancel diving operation for today (see Deck Cam feed). We will complete our CDT casts and continue with sonar surveys.

We will remain at Barkley Canyon in the hope that the weather abates tomorrow morning.


May 20th 14:30 Update

20-May-13

With the successful deployment of the Folger Deep platform, our inshore engineering work is now complete. We are heading offshore to our Barkley Canyon sites to prepare for ROV dives tomorrow morning. On our way to Barkley we will conduct CTD casts and then start a new series of sonar surveys in a location suspected of having gas seeps. Our dives should start at 0730 Tuesday morning, weather permitting.

0730 - Dive at Barkley Hydrates site, sonar tripod positioning and visual survey

1330 - Dive to location of gas seeps detected during sonar survey


May 19th 11:50PDT Update

19-May-13

We have just successfully recovered the VPS platform. The operation was completed very smoothly by the Tully deck crew. The sun is shining out here, but the seas have been building, so the captain has made the decision that we should not dive this afternoon.

Instead, we are going to do some more deep water CTD rosettes and continue the echosounder survey for bubble plumes from last night. The current plan is still to deploy Folger Deep Monday morning.


May 19th 0050PDT Update

18-May-13

We're right on schedule at the time of this update:

Overnight: Echosounder survey (from Tully)

Sunday morning: Dive to recover VPS platform

Sunday afternoon:

  • Barkley Hydrates
    • Deployment of second sonar tripod on winch
    • Dive for visual survey of Wallyland; positioning of tripod; coral cliff survey
  • Echosounder survey (from Tully)


May 18th 0800PDT Update

18-May-13

Here is an update of our operations and plans for Saturday and Sunday:

  • CTD rosettes in Effingham Inlet
  • Science dives in Effingham Inlet with Verena Tunnicliffe
  • Redeployment of Folger Deep
  • Echosounder surveys looking for bubble plumes
  • Potential dives at identified locations.

 


May 17th 1630PDT Update

17-May-13

We are in Barkley Sound, waiting for a crew change to be completed.

  • Ian Kulin, Jonathon Miller, Reece Hasanen and Reyna Jenkyns are disembarking.
  • Adrian Round (as expedition leader), Ian Beliveau and Karen Douglas are joining us.
  • Verena Tunnicliffe (UVic) and Bob McDonald (of CBC's Quirks & Quarks) are spending one night with us during our exploration of Effingham Inlet.

Here is an update of our operations and the plans for the next few days:

Currently:

  • Tully in Barkley Sound; crew change in progress from Bamfield
  • Folger Deep platform on board for maintenance
  • PODs 1, 2, 3, 4, and Barkley Hydrates have been redeployed.

Today and tomorrow:

  • CTD rosettes in Effingham Inlet
  • Science dives in Effingham Inlet
  • Redeployment of Folger Deep


May 17th 0800PDT Update

17-May-13

Currently staging Oceanic Explorer dive OE0060 at Folger Passage. Goals for this dive:

  • Recover hydrophone
  • Disconnect bottom pressure recorder
  • Disconnect Folger Deep instrument platform
  • Recover existing cable to node
  • Prepare Folger Deep instrument platform for recovery to ship


May 15th 2000PDT Update

15-May-13

Dive OE0057 was completed with an early end due to ROV issue.

  • POD 4 reconnected (although not powered because Barkley Hydrates is not yet connected).
  • Unfinished at POD 4 due to early end: CTD on tripod not in place.
  • POD 3 recovered; will be maintained overnight
  • No overnight dive.


May 14th 2130PDT Update

14-May-13

Today, we:

  • Deployed Pod 4
  • Dove to Pod 4 (dive OE56), but ROV lost power and dive was aborted 
  • Deployed one tripod for new sonar installation at Barkley Hydrates via winch
  • Completed maintenance on Barkley Hydrates platform
  • Deployed Barkley Hydrates platform via winch

Tomorrow's plans are to:

  • Recover Plankton Pump (Dive OE57 beginning at ~7:30 am)
  • Connect Pod 4 (Junction Box 05)
  • Deploy CTD tripod 
  • Connect Barkley Hydrates instrument platform
  • Tasks for Kongsberg sonars
    • Locate and inspect winch-deployed tripod 
    • Identify landing zone for second tripod 
    • Survey proposed deployment sites
  • Wally Land tasks 
    • Inspect shell and wood experiments 
    • Recover yellow markers 
    • Perform photomosaic survey
  • Place marker beacon at Pod 3
  • Recover two parking positions
  • Recover Pod 3


May 14th 08:00PDT Update

14-May-13

Barkley Canyon Pod 4 is being deployed from ship. Once deployment is finished, the Oceanic Explorer will be launched for dive OE0056 to connect Pod 4 and prepare Pod 3 for recovery.

Barkley Hydrates Instrument Platform remains on deck for now.


May 13th 18:40PDT Update

13-May-13

Dive OE0055 complete. ROV on deck.


May 13th 14:13PDT Update

13-May-13

Dive OE0055 is on. Diving to Barkely coral cliffs.

Location: Barkley Canyon                                                                        

Objectives

·  Deploy autonomous Nortek profilers (DeviceIDs: 23163,23164)

·  Visual survey of coral cliffs

Full Dive Plan OE0055

 


May 13th 12:40PDT Update

13-May-13

Dive OE0054 is complete. ROV on deck; the Barkley hydrates platform is coming up on board first time since 2009. 


May 13th 08:55PDT Update

13-May-13

Dive OE0054 is on.

Location: Barkely Canyon

Objectives:

·  Place marker beacon

·  Disconnect Wally

·  Disconnect Barkley Hydrates Instrument Pod and clear cables safely away

·  Recover Barkley Hydrates Instrument Pod (DeviceID: 10002)

Full Dive Plan OE0054

 


May 12th 20:15PDT Update

12-May-13

Dive OE0053 is complete. ROV and Instrument Pod (POD) 4 on deck.


May 12th 10:15PDT Update

12-May-13

Dive EO0053 in progress.  

Location: Barkley Canyon

Objectives:

· Connect POD 1 (DeviceID: 10011) to network

· Deploy Dragonfish camera system (DeviceID: 23152, 23074)

· Deploy Nortek Profiler (DeviceID: 11302)

· Deploy Kongsberg (DeviceID: 11301)

· Visual transect survey

· Organism sampling

· Recover bait trap

· Recover marker beacon

Full Dive Plan OE0053


May 12th 06:12PDT Update

12-May-13

Dive OE0052 is complete. ROV on deck.


May 11th 18:45PDT Update

11-May-13

Dive OE0052 is on. Weather remains to be a constraint.

Location: Barkley Canyon 

Objectives (updated @ 19:30):

·         Prepare POD 4 Recovery

o   Visual inspection

o   Visual transect survey

o   Prepare platform for recovery

o   Disconnect POD 4 from network and clear cables safely away

·         Prepare POD 3 Recovery

o   Visual inspection at POD 3

o   Prepare Kongsberg (DeviceID: 22789) for recovery

o   Recover Nortek Profiler (DeviceID: 12003) and secure to IP

o   Recover Sediment Trap (DeviceID: 12005) and secure to IP

o   Disconnect POD 3 from network and clear cables safely away

·         Place marker beacon at POD 1

·         Visual survey from POD 1 to Hydrates IP

·         Inspect shell/wood experiment

·         Survey area around Barkley Hydrates for later maintenance work.

·         Recover POD 4

 

Full Dive Plan OE0052


May 11th 16:30PDT Update

11-May-13

Dive OE0051 Complete.


May 11th 11:15PDT Update

11-May-13

Dive OE0051 is on. ROV is in the water. Live video is streaming.


May 11th 08:00PDT Update

11-May-13

Preparing for dive OE0051. Weather is a constraint today.

Location: Barkely Canyon

Objectives:

·   Connect POD 2 JB-03 (DeviceID: 10018) to network

·   Deploy Sediment Trap (DeviceID: 11008)

·   Deploy Dragonfish camera system (DeviceID: 23153, 23076)

·   Deploy Nortek Profiler (DeviceID: 11203)

·   Deploy Kongsberg (DeviceID: 12007)

·   Visual transect survey

·   Organism sampling

·   Recover bait trap

 

 

Full Dive Plan OE0051

 


May 10th 14:00PDT Update

10-May-13

Preparing for the next Dive – OE0050. Scheduled to start early PM.

Location: Barkley Canyon 

Objectives:

 

·  Secure floats to camera tripod

·  Visual inspection of Instrument Pod 

·  Deploy bait trap

·  Visual star-configuration transect

·  Prepare Kongsberg (DeviceID: 11301) for recovery

·  Recover Nortek Profiler (DeviceID: 11302) and secure to IP

·  Recover camera system (DeviceID: 12170, 12126) and secure to IP

·  Disconnect POD 1 (DeviceID: 10011) from network and clear cables safely away.

 

Full Dive Plan OE0050

For details, see Daily Log 10 May


May 10th 07:00PDT Update

10-May-13

Dive OE0049 dive is scheduled for 7:30am. The plan is to complete remaining tasks from OE0048 dive and replace the CTD for Upper Slope. 

Location: Barkley Canyon

Objectives:

• Deploy CTD tripod (DeviceIDs: 23031, 23047) at Upper Slope

• Visual Upper Slope IP (DeviceID: 10003) Inspection • Recover CTD (DeviceID:22607) at Upper Slope

• Recover Dragonfish camera system (DeviceID: 23072, 12127) and secure to Instrument Pod

• Recover Sediment Trap (DeviceID: 11008) and secure to Instrument Pod

• Inspect VPS Platform (DeviceID: 22612)

• Disconnect POD 2 from network and clear cables safely away. 

Full Dive Plan OE0049


May 9th 14:55PDT Update

09-May-13

Planning the first dive of the Leg 2 - Servicing NEPTUNE observatory. Dive OE0048. 

Location: Barkley Canyon

Dive  Objectives: 

• Visual IP Inspections

• Position marker beacon

• Deploy bait trap

• Visual transect

• Prepare Kongsberg (DeviceID: 12007) for recovery

• Recover Nortek Profiler (DeviceID: 11203) and secure to IP

• Recover Dragonfish camera system (DeviceID: 23072, 12127) and secure to IP

• Recover Sediment Trap (DeviceID: 11008) and secure to IP

• Inspect Vertical Profiler System (VPS) Platform (DeviceID:22612)

• Disconnect POD 2 from network and clear cables safely away

• Recover CTD (DeviceID:22607) at Upper Slope.

Full Dive OE0048 Plan

For details, see Daily Log, 9 May  


May 8th 22:30PDT Update

09-May-13

Leg 2 (May 9-28) of the "Wiring the Abyss 2013" on Tully begins.

The plan for the second leg of the “Wiring the Abyss 2013” aboard CCGS Tully is to service instruments at two locations of the NEPTUNE observatory of Ocean Networks Canada. In particular we will focus on Folger and Barkley Canyon locations off the west coast of Vancouver Island.

The planned service includes:

·  Recover, perform maintenance and redeploy Folger Deep platform

·  Recovery, maintenance and redeployment of four Barkley Canyon pods (instrument platforms) and surrounding instruments

·  Recover and perform maintenance and deploy at Barkley Hydrates

·  Install sonar system at Barkley Hydrates

·  Perform sampling at all Folger and Barkley sites and in Effingham Inlet.

For details, see Daily Log, 8 May


May 8th 13:45PDT Update

08-May-13

The Tully has returned to IOS in Patricia Bay for a science crew exchange and re-mobilization for the second leg of this operations and maintenance expedition. Plans for today include the installation of a new satellite system through Oceaneering. Tomorrow, Leg 2 of the cruise should see us heading out into the Northeast Pacific Ocean to service NEPTUNE observatory sites Barkley Canyon and Folger Passage. Taking the reins from Richard Dewey, we will now be hearing from Expedition Leader Ian Kulin, who is Ocean Networks Canada's Associate Director for Marine Operations.


May 6th 23:30PDT Update

07-May-13

Tomorrow (May 7), Verena Tunnicliffe will join us to complete a two-year research effort. Planned are video survey dives up into Patricia Bay, and at least two vertical video profiles along the steep cliff walls deeper in the Inlet. And then Leg #1 will come to an end on May 8. Leg #2 of the Tully cruise will be devoted to servicing several off-shore sites of the NEPTUNE observatory. 

For details, see Daily Log, 6 May


May 5th 23:50PDT Update

05-May-13

The ship is now steaming towards Saanich Inlet. Over the next few days we will re-assemble the Saanich Inlet Instrument Platform and the Oregon State University-led Benthic Oxygen Flux experiments. Benthic surveys are required for two camera projects, and during the last dive we will recover a camera for maintenance. All going well, we should be back at the Institute of Ocean Science dock for the exchange by early afternoon, May 8.

For details, see Daily Log, 5 May 


May 5th 09:45PDT Update

05-May-13

Currently preparing for a dedicated dive to investigate and sort out the navigation system.


May 5th 00:05PDT Update

04-May-13

Yesterday, the Bottom Boundary Layer platform was deployed and connected to the Strait of Georgia Central Node. Today's plans call for deployment of the Seismic Liquefaction In-situ Penetrometer (SLIP) and Delta Hydrophone Array near the Fraser Delta node, followed by possible deployment of the Eastern Hydrophone Array near the Strait of Georgia Eastern Node.

For details, see Daily Log, 4 May


May 4th 00:20PDT Update

03-May-13

Today we will finish the sonar surveys at 6AM, and steam to our SoG Central Node site for one platform deployment (the Bottom Boundary Layer package) and one platform recovery (the autonomous CTD mooring). During these deck operations, we will assemble the complicated Seismic Liquefaction In-situ Penetrometer (SLIP). It will be deployed during a slack tide in the early afternoon. This leaves most of the day for Tom to re-assess his hydrophone arrays, and deploy them on Sunday.

For details, see the Daily Log, 3 May


May 2nd 22:00PDT Update

03-May-13

Tomorrow we pick up two of Tom Dakin's students (Hannan and Kristen) and Gwyn Lintern of NRCan from Steveston, and continue the deployments near the Delta Node. First up will be the Delta Dynamics Laboratory (DDL) at the end of its own kilometre extension cable, up-slope from the Delta Node. This cable is presently being spooled onto the ROV reel. We will now add an extra dive at the end of the day to recover the ONCCEE platform.

For details, see the Daily Log, 2 May


May 1st 23:00PDT Update

02-May-13

The ship slowly streamed into the Strait of Georgia early this morning. After a few days of strong winds, we brought calm to the Strait. But the mighty Fraser is cranking up it's flow, as the Rocky Mountain snow begins to melt. With a 5 fold increase in fresh water volume in May, there is an equal increase in the suspended sediment. The plume had been detected in our Ferry data to have been broadening over the last few weeks, and it clearly now fills a majority of the southern Strait.

The next few days we be at the Delta Laboratory site. After a staff picking up at Steveston early May 2, we will scout out the northern end of the Delta cable, and get the Delta mini-Node re-built, deployed, and connected. The first of two tech-demo hydrophone deployments may finish the day. On May 3, Gwyn Lintern (NRCan) will join us and we will lay a cable and deploy a new DDL platform. May 4 will focus our efforts on the Seismic Liquefaction In-situ Penetrometer (SLIP). This leaves a final hydrophone array and the Bottom Boundary Layer (Dalhousie and MUN) platform for SoG Central.

For details, see the Daily Log, 1 May


May 1st 08:00 Update

01-May-13

Today we'll focus on the recently (April 26) laid Delta extension cable. Our first dive will connect the south end to the SoG East Node. Another dive at the north end of the extension cable will be necessary to secure it, and determine where to deploy the new Delta mini-Node. Time permitting, we'll deploy and connect the mini-Node. The Fraser River freshet is starting, so we might encounter strong surface currents, laden with silt and sediments.

First Dive, diving soon:

Dive OE0028 Objectives:

  • Find Node and East VIP
  • Head east and find cable
  • Follow south and get cable end, take to Node, connect
  • Head SSE and locate autonomous platform, recover

 


April 30th 23:00 Update

30-Apr-13

We finished rigging the navigation beacon assembly and got away from the dock reasonably early. A multitasking effort to test the ROV navigation system while conducting the first CTD and Rosette cast, accommodated a first dive before lunch. This went smoothly. The afternoon saw another dive and platform recovery and the deployment of a vital autonomous mooring at the entrance to Saanich Inlet. We deployed the new forensics platform after dinner and then dove to connect it. A quick return to the IOS dock to exchange a dozen platforms, and we're now heading over to the Strait of Georgia.

For details, see the Daily Log, 30 April.


April 29th 20:00 Update

29-Apr-13

We will either sail at 06:15 or 07:15 tomorrow morning, depending on status of ROV systems. Plans for tomorrow include:

  • CTD and Rosette at SI VIP site/depth to calibrate oxygen measurements.
  • First dive to recover bones from pigs 7 and 8 and HD webcam-platform. Transfer all gear to forensics platform.
  • Second dive to recover SI VIP.
  • Deploy new forensics platform with pigs 9 and 10.
  • Deploy ROV to connect forensics.

On board today for forensics operations: Dr. Gail Anderson and Chantal Turpin

Report by Dr. Richard Dewey, Expedition Leader

For details, see the Daily Log, 29 April


Change to Twitter

29-Apr-13

On 29 April 2013, our Twitter handle for expedition operations changed from @nc_operations to @oceanetworksops.


Cruise Season Begins

29-Apr-13

On April 29, ONC embarks on a series of maintenance cruises to service both the VENUS and NEPTUNE cabled observatories. Over the next two months, we’ll execute three of our eight 2013 expeditions, including two back-to-back cruises on the CCGS John P. Tully (April-May), servicing facilities first in the Salish Sea (VENUS) and then off-shore at Folger Passage and Barkley Canyon (NEPTUNE). 


Fall Cruise Canceled

25-Sep-12

Our efforts to find a suitable replacement vessel for the T.G. Thompson have come to an unsuccessful end and we are now obliged to cancel our Wiring the Abyss 2012 fall maintenance cruise, originally scheduled for September 24 to October 10. We are now beginning to plan a 3-week cruise in April/May 2013 during which we would complete work from this fall in addition to other maintenance work that may arise over the winter.

Thanks to all for their patience while we searched for a solution to the loss of the Thompson, which is out of action for several months for repairs to its starboard drive and steering mechanism. Everyone receiving this message will be advised in early January when have firm dates for our next cruise.


Fall Cruise Ship Still Uncertain

19-Sep-12
The choice of possible vessels for the fall cruise has been narrowed down to one candidate. The feasibility of mobilizing ROPOS on that vessel is now under investigation. Based on the ship’s availability and mobilization time required, the earliest a cruise could occur is the beginning of October. There is a second vessel under consideration. he same issues regarding mobilization of ROPOS are under investigation for that ship. As these investigations are still ongoing, the final decision (cruise versus no cruise) has been postponed to next week. A further update will be issued on September 24.


Status of Fall Cruise

14-Sep-12
We are still searching for a solution to our support vessel problem. Two candidate replacement vessels are still being considered and discussions are underway with their operators. We hope to reach a final decision (cruise versus no cruise) early next week. A further update will be issue on September 19.


June 25th 2300 Update

25-Jun-12
The last day of Wiring the Abyss 2012 was one of the best of the cruise. The weather offshore was beautiful: light winds at 10 knots, calm seas and bright sun. We could see the mountains of Vancouver Island all the way from Barkley Upper Slope where we did our final operations of the cruise.

To start the day, Dive 1573 began around midnight to redeploy the repaired camera at POD#4 (mid-canyon wall). The camera was found to have connector issues with its pan and tilt system on the first leg of the cruise, so it was brought on deck and creatively repaired using components adapted from other systems. The camera worked perfectly on deployment. The morning was spent doing CTD casts and water sampling to facilitate calibration of the Vertical Profiler System (VPS) and its onboard instruments. On the last dive of the cruise, Dive 1574, we swapped the Upper Slope Bottom Pressure Recorder (BPR) for a new one and then spent some time observing the VPS performance underwater with ROPOS. It was very exciting to see an instrument connected to the seafloor network at 400m operating in the photic zone!

We are now on the way back to the Esquimalt Graving dock, arriving home at 0800 Tuesday, June 26th, after a busy and successful month at sea.


June 24th 2030 Update

24-Jun-12
VPS Success!

This afternoon, in bright sun, very calm winds and a long period 1m swell, we decided that conditions were optimal for a Vertical Profiler System (VPS) surfacing test. The system was first brought to 10 and then 5 metres, while we observed cable tension and the behaviour of the winch drum and sheave. At 5 metres, there were occasional zero tension spikes on the main VPS cable but no worrisome behaviour at the seafloor winch. We switched on the radiometers at this point. We then made the decision to slowly bring the system up to 1 metre.

We put a small boat in the water and set up a video camera on the bow of the Thompson. As the ship approached the VPS location, we were able to see the float even at 3 metres depth. We repositioned the ship, and the zodiac to be as close as possible to the VPS which we then raised and held at 1m depth. At this depth, the VPS broke the surface several times, without incident.

We are now conducting calibration tests with the Thompson CTD-rosette. If there is sufficient time tomorrow morning, we will put ROPOS in the water to observe the VPS docking and undocking, and then film it on the way back up to the surface.


June 24th 1400 Update

24-Jun-12
We have completed operations at ODP 1027, and we are now back at Barkley Canyon. It's a beautiful sunny day here, and we already spotted a shark and a two whales from the deck. Early this morning, POD#4 was recovered for installation of the newly refurbished camera and pan & tilt system. While that work is happening on deck, we transitted to Barkley Upper Slope where the CTD casts for the Vertical Profiler System (VPS) calibration and VPS testing are underway.


June 22nd 0645 Update

22-Jun-12
After a long day of careful preparations on deck by ROPOS and the engineering team, ROCLS (Remotely Operated Cable Laying System) was ready to be lowered to the seafloor on ship's winch near the ODP 1027 instrument platform with the 20km northeast tsunami array cable aboard. ROPOS started the descent for Dive 1570 around 2115 PDT (0415 UTC). The first tasks on this dive were to finish some operations from Dive 1557 which were not completed due to problems with one ROPOS manipulator arm which caused that dive to be aborted early. Following that, a test was performed on the cable before the lay to ensure that no faults were introduced during its the trip to the abyss. The cable is intact, and so the cable lay was started. We estimate that this lay will take around 40 hours, after which we will connect the BPR which is being carried out with ROPOS and ROCLS to the end of the cable line. Last night, the crew and science party enjoyed a barbeque dinner on deck complete with a fabulous visit from dozens of Pacific white-sided dolphins surrounding the R/V Thompson.


June 21st 0430 Update

21-Jun-12
The 3.5km cable originating at the Endeavour node and extending halfway to the Mothra instrument platform has been laid. As we neared the ridge, the terrain became rougher with some deep crevasses and ridges along the route. A visual inspection of the rough area of the lay did not reveal any noticeable problems with the cable, so we are happy with the results of Dive 1569. We are now en route to ODP 1027 to lay the second cable in the tsunami meter. This 20km cable will connect the northeast Bottom Pressure Recorder (BPR NE) to the ODP 1027 instrument platform. Preparation, cable lay and testing is anticipated to take several days at the site. The weather is now very good for operations - wind speed 11 knots and calm seas.


June 20th 1700 Update

20-Jun-12
We are in the middle of our last dive of this cruise at Endeavour. ROPOS is on the bottom surveying the northern half of the cable route connecting the Endeavour node to the Mothra instrument platform (IP). Earlier in this cruise, we connected two Regional Circulation Moorings (RCM) to the Mothra IP, the southwest and southeast. At this point, all four planned moorings are deployed, but the two southern ones have yet to be connected to the backbone of the network. In this cruise, we are only able to lay the first 3.5km of the Mothra cable, but this sets the stage to connect the south moorings on the fall cruise when we will complete the cable lay. Yesterday and into the early morning today, we successfully finished the recovery of the instruments from Main Endeavour Field. Tempo-Mini is now on deck and cleaned up and the Benthic and Resistivty Sensors (BARS) was also brought on board. A highlight yesterday was the release of the Remote Access water Sampler (RAS). It was tethered with a pin securing it above a diffuse flow vent on Grotto and held upright by a float. After ROPOS pulled the pin, we all stood on deck in the sunshine trying to be the first to spot it at the surface. Following the cable lay now in progress at Endeavour, we will transit back to ODP 1027 where we will continue installation of the tsunami meter.


June 19th 0320 Update

19-Jun-12
We are currently diving at Main Endeavour Field. The wind died down this afternoon, and then it took a few more hours of patient waiting for the sea state to come down enough for us to dive. Dive 1567 started by successfully moving and connecting the Northeast Regional Circulation Mooring (RCM NE). We then transitted underwater to Main Endeavour Field which took just over 3 hours, and we are now preparing some instruments for recovery on a later dive. COVIS (Cabled Observatory Vent Imaging Sonar) appears to be in very good condition, and will remain in operation. Tempo-mini, an integrated suite of instruments for studying vent communities, BARS (Benthic and Resistivity Sensors), a probe for measuring properties vent fluid emissions and RAS (Remote Access water Sampler) will be recovered this cruise. So far, the BARS cable was found to be melted through which certainly explains its failure to convey data for the last few months. The connector to RAS was completely covered by growth: a community of tubeworms moved in. Further updates to follow!


June 18th 1040 Update

18-Jun-12
Waiting on the high seas. Nothing new to report; wind is still at 20 knots and seas are too high for diving. We are working on documentation and exploring the thought problem of how to build an accurate weighscale on the moving ship.


June 17th 0645 Update

17-Jun-12
The weather co-operated, and we were able to start Dive 1566 at around 2030 PDT. We are just over halfway through laying the cable to connect the RCN END IP to the node. When this cable is down, we will bring the short period seismometer and BPR online. We will also reconnect the RCM NW mooring on this dive. On the next dive, we will move the RCM NE mooring into position and connect it. Following that, we're heading to Main Endeavour Field for some eagerly anticipated maintenance.


June 17th 2015 Update

17-Jun-12
WEATHER NOTICE - Dive 1567 RCM North and Main Endeavour Field scrubbed due to weather. Forecast for building wind and seas tonight. We will re-evaluate at first light.


June 16th 0046 Update

16-Jun-12
It has been a very busy few days for the crews of the R/V Thompson, the CCGS Tully, ROPOS and the the NEPTUNE team. Prevailing wind direction and the satellite receiver being in the shadow of some critical structure on the boat has left us with the challenge of a spotty internet connection. We moved RC-South IP to its position selected on the previous dive R1563. ROPOS connected RCM Southeast Mooring to RC-South IP and laid the cable between RC-South IP and RCM-SW. The Cable route from RC-South IP to RCM-SW mooring was surveyed and the best location for RCM Southwest mooring was selected. We took the opportunity to visually navigate the RCM Southeast and RCM Southwest Moorings to their respective deployment sites, connected them and inspected the moorings. Beginning early and working in concert with the CCGS Tully we released the RCM-NE mooring and recovered RCN END IP. ROPOS also moved over to BAMBI (High Rise vent field) to pick up BARS and then returning to assist in the RCM-NE deployment. We also surveyed RCM Northwest mooring over its full length to check the condition of connections.


June 16th 1100 Update

16-Jun-12
We are waiting on the weather! The wind has picked up to 20-25 knots, and the swell is increasing. Unfortunately because the wind is coming from the SW, our heading has also been SW for most of the last 24 hours. At that heading, we sometimes lose sight of the satellite, which is the reason for our spotty communications over the last 2 days.

We are all ready for Dive 1566 in which we will lay cable to connect the RCM NE mooring that was deployed yesterday with the Tully. Due to the proximity of the RCN END IP target position and the final position of the mooring, we had to place the mooring 250m away from its eventual site to avoid the danger of two taught cables in the water at the same location when ROPOS descends with the IP. Once the IP is in place, we will connect the short period seismometer, BPR and RCM NW moorings which are already at the site, and then lay the cable to the Endeavour node.

On the following dive, we will reposition and connect the RCM NE mooring. However, until the weather allows us to start diving again, we are spending our time doing multibeam surveys and CTD casts in the Endeavour region.


June 14th 0825 Update

14-Jun-12
Dive 1564 started around 0030PDT, and is still in progress. On this dive, we are coordinating with the CCGS John P. Tully to deploy two regional circulation moorings (RCM) at Endeavour: RCM Southwest and RCM Southeast. The dive began by repositioning the RCM South IP and laying the cable between the IP and the SW mooring site. It is a cool grey morning out here at 8 deg C, and the Tully is visible in the mist off the stern of the Thompson. This dive will continue through the daylight hours until ROPOS is recovered this evening. Next we will proceed to RCM NE where the RCM NE mooring will be released by ROPOS and recovered by the Tully tomorrow morning. A replacement mooring will be deployed at that location. On another note, tests with the VPS are going well. The motor issues from earlier in the cruise have been resolved, and the VPS is performing monitored profiles of the water column at this time.


June 13th 2000 Update

13-Jun-12
Dive R1653 is complete. Completed objectives were to lay cable from Mothra Junction box to the RCM South instrument platform. Special equipment will be required on subsequent dive to lift and move the RCM-South instrument platform to a more accessible location. A new location for RCM-South Instrument platform was selected, and a location for the RCM-SE mooring which will be deployed tomorrow AM. A landing site was also selected for ROCLS for the next cable deploy from RCM-SE to RCM-SW. Two sediment cores were collected for Laurence Coogan.


June 13th 0845 Update

13-Jun-12
The transit to Endeavour Mothra was a little bumpy, but we made it on schedule. Dive 1562 completed early this morning with a successful deployment of the Mothra IP. It took several attempts to find a site which was solid enough so that the platform was level and not sinking in the sediment. Finally, a good site was found, which is in reach of the future BARS and short period seismometer deployment sites. Several push cores and a sulfide sample were taken before ROPOS was recovered to the surface. The ROPOS control room was full of tourists enjoying the spectacular black smokers at Mothra. The wind has died down and the seas are calm enough to start Dive 1563, in which the cable between the Mothra IP and RCM South IP will be laid.


June 12th 0200 Update

12-Jun-12

The first day of leg 2 of the cruise began with a crew change: 12 members of the science party disembarked and 11 joined the cruise. Once the members were onboard the R/V Thompson, operations began to retrieve the Folger Deep instrument platform. The platform has has been cleaned, serviced and tested on deck and is ready to be redeployed with a new high frequency hydrophone and ACDP installed. Dive 1561 to redeploy the Folger Deep IP will be starting shortly.  Following that, the pilot will be picked up by the Barkley Star in the early morning, and we will transit to Endeavour to begin operations at Mothra.


June 12th 0845 Update

12-Jun-12
Dive 1561 completed at approximately 1340UTC (0640PDT). The Folger Deep IP was successfully redeployed after cleaning and installation of the new high frequency hydrophone and ADCP. Communications with the instruments was tested, and initially all were operational. After a few minutes, the CTD showed evidence of a ground fault. A suspected driver issue caused communication problems with the junction box, so the CTD could not be restarted. These problems are under investigation. We picked up the Folger Pinnacle cable that was found to be faulty and recovered ROPOS. We are currently underway to Endeavour to start by deploying the Mothra IP. Sea state is calm, wind speed 10 knots.


June 10th 2000 Update

10-Jun-12

The day closed with a big challenge facing the Wiring the Abyss 2012 team – after re-laying the repaired 23 km long tsunami sensor cable, we found that it had been damaged, once again, during this current deployment. The reasons for the damage are not yet clear. We are currently heading to Bamfield for the scheduled science party crew change. This will give the engineering team time to evaluate the procedures, identify the fault, and propose a solution. While in the Bamfield area tomorrow, repairs at the coastal node are of Folger will take place.


June 10th Update

10-Jun-12

Termination of the newly cut end of the now 23 km long tsunami array cable (the first of three) was completed early this morning. Following a detailed and careful deck deployment strategy, the termination canister with the long cable trailing was deployed on ROPOS and is now on its way to the bottom. Once there the cable will be tested and connected to the bottom pressure sensor. ROPOS will be at work on the bottom by early afternoon. So, stay connected for news about the tsunami sensor connection!


June 9th 1800 Update

09-Jun-12

Wiring the Abyss 2012 is still on location at the end of the newly laid tsunami sensor array cable at NEPTUNE's ODP 1027 site. Good progress is being made on terminating the connector end of the cable that was recovered yesterday. We anticipate that this work will be completed very early tomorrow morning around day break. So, this evening, once again we remain in place, anchored to the end of the now 23 km-long cable. In the meantime, preparations are underway to deploy the terminated end of the cable and connect the bottom pressure sensor to it on the seafloor. After this operation is completed, we'll transit to Bamfield for a science party change on Monday morning. After the crew change, the ship will head to the Folger area.


June 8th 2300 Update

08-Jun-12

World Oceans Day was chalk-a-block full of on deck action onboard the R/V Thompson. The damaged end of the first tsunami array cable was brought up to the sea surface by ROPOS where a team of NEPTUNE and ship's officers and crew were waiting for its hand-off. After ROPOS transferred the cable end with its termination can to floats, it was moved through the water along the edge of the ship to the stern beneath the ship's A-frame. From there, the cable end was secured to a line connected to a motorized capstan, which upon turning, began bringing the cable onboard the ship. Because there was cable damage at two locations: the very end of the cable and at 2 km from the end, we had to bring back aboard and cut off a whopping 2 km length of cable. This required most of NEPTUNE's science team to complete this task because the cable had to be spooled onto a drum on deck using human-power. Everyone on shift took turns turning the spool and the job was completed by early evening. After securing and protecting the cable on deck, the detailed task of building a new termination connector for the cable began. So now we sit on this location for the next day or so, anchored to the seafloor by the end of a now 23 km-long cable, standing vigil until the new connector is complete. Once completed, the cable end will be lowered back to the seafloor with ROPOS, connected to NEPTUNE's network, tested, and finally reconnected to the first of three new tsunami wave sensor instruments.


June 8th 1130 Update

08-Jun-12

World Oceans Day began onboard the Wiring the Abyss cruise with very positive vibes and a bit of a buzz. After the successful test of the cable yesterday where we found that the damage to the first of three tsunami array cables was isolated at one end (yeah!), the ROPOS and NEPTUNE teams along with the terrific R/V Thompson officers and crew worked early this morning to develop and now implement a plan to recover the far end of the cable (where the bottom pressure recorder is located), secure it on deck, cut off the bad bits, and then establish a clean environment close to the cable end for terminating this newly cut end. Whew, it makes me tired just writing about it, but this group of experienced and capable folks not only know how to plan it, they can do it too. So, for the next almost two days, we'll be conducting this complex operation. As before, we'll post updates on a regular basis, so stay connected.


June 7th 0100 Update

07-Jun-12
This afternoon was a landmark for ROPOS where they beat the record for their longest dive to date, three days. The dive was epic because of the challenging 25 km long cable lay – the first of three cables that make up the new tsunami sensor array. The cable lay was completed last night at 1800 hours and the equipment, including the bottom pressure recorder that senses the surface tsunami wave, was connected. Currently the fully laid and connected system is going through de-bugging and testing because the instruments are not yet talking with NEPTUNE's network. For the first test ROPOS will use an optical time-domain reflectometer, or OTDR for short, to pulse light down the 25 km cable and measure the light reflected back. With the results, we will be able to assess the integrity of the cable. The weather is expected to get worse later this morning, so ROPOS will be recovered after the test for some well-deserved maintenance. In the meantime, we will conduct some multibeam surveys in the area of IODP 1027.


June 7th 2000 Update

07-Jun-12

Good news resulted from the optical time-domain reflectometry testing of the recently laid tsunami array cable. The cable fibres are good from the junction box out to 23 km, so after the rest of the maintenance work is completed on this current dive, ROPOS will return to the ship for a 23 km transit to repair the end of the cable. Preparations are now underway to bring the end of the cable onboard so that the damaged end can be removed and the new end re-terminated.


June 7th 1300 Update

07-Jun-12

As we wrap up the multibeam survey to fill in bathymetry data gaps in the 1027 area, the ROPOS and NEPTUNE engineering teams are prepping for the next dive. The first priority will be further testing of the 25 km tsunami array cable that was laid over the past few days. Today's early AM optical time-domain reflectometer (OTDR) test showed that the cable had a defect, likely at the terminus end. For the next OTDR test, we will shoot the light from the other end of the cable to further narrow down the location of the defect. Once identified, we'll bring the cable on deck and repair it. On this dive, ROPOS will also swap-out an auxiliary platform, which includes an aquadop current meter and a differential pressure gauge for maintenance.


June 5th 1715 Update

05-Jun-12

The afternoon's cable laying experienced a small bump in the road when the chain that controls the spooling of the cable broke. The experienced ROPOS team ably stepped up to this challenge by removing the chain and quickly establishing a new cable laying protocol in tandem with the ship so that the cable could be laid with the spool in a freewheel mode. With this new approach in place, it will take longer to complete the first cable lay. So, for the foreseeable future, we'll just be cabling and cabling along.


June 5th 1400 Update

05-Jun-12

Overnight, ROPOS continued to lay the first of the three tsunami cables. With the exception of a short bit of down-time this morning to correct some pesky stray cables, the laying procedure continued this morning and half of the 25-km length was deployed by noon.


June 4th 1930 Update

04-Jun-12

Since arriving at site 1027, preparations have been underway to lay the tsunami sensor array cables. After lowering the 25 km long cable spool to the seafloor, ROPOS dove to the sea bottom to set up this highly technical and challenging operation including set-up of a seafloor staging area 2.6 km below the sea surface, positioning the cable connectors, establishing a robust communication system with the ship, which has to move in tandem with ROPOS, and confirming and re-confirming a safe route. With the set-up complete, cable laying began early this evening and we anticipate the first cable will be finished by this time tomorrow. After the first cable is laid, we will repeat the process for the second of three tsunami sensor array cables. The third cable will be installed this fall to complete the array.


June 3rd 2320 Update

03-Jun-12

It's almost midnight and we have just arrived at our next operations site – ODP 1027, the tsunami sensor array location. During the transit here, the ROPOS and NEPTUNE teams prepared the vehicle, cables, and instruments to begin the installation of the array. The first step is to lay a 25 km long cable that will connect one of three bottom pressure recorders to NEPTUNE's power and internet cable system. Connecting these pressure sensors to the internet forms the backbone of the tsunami wave sensing system and this cable is the first of three that make up the sensor array. This first cable lay and pressure sensor connection will be completed over the next two days.


June 3rd 1430 Update

03-Jun-12

ROPOS is currently on its way back to the ship after an exciting and more than 12-hour long dive. The auxiliary broad brand seismometer platform at 889 was replaced, a new sonar was installed at a locale rife with bubbles coming out of the seafloor, and an autonomous osmo-sampler was recovered after more than a year collecting sub-seabed water samples. The seafloor in this area is quite dynamic – gas bubbles with a hydrate coating are released in areas that change with time. An area that was active in 2010, dubbed "bubbly gulch" seemed relatively inactive today. Yet another area was extremely active with gas bubbles streaming out of the seafloor at a rapid pace that attracted crabs. The crabs appeared to be dining on the bubbles and enjoying a major picnic! We also visited the carcass of a whale that had been discovered earlier by scientists from the Monterey Bay Aquarium Research Institute and found only very friable bones left on the seafloor.

We're now wrapping up work here at 889 and after a CTD cast, we'll get underway and head to 1027 to install the first of three new bottom pressure sensor that are part of the new tsunami sensor array. This work should commence around midnight tonight.


June 2nd 2300 Update

02-Jun-12

The first dive at 889 was completed this evening with quite positive results. ROPOS dove to the Integrated Ocean Drilling Program borehole 1364A, which is installed with an advanced circulation obviation retrofit kit (ACORK). The ACORK 1364A extends 300 m below the seafloor and is equipped with pressure gauges that monitor fluid pressures at sub-seabed depths of 155, 205, 245 and 295 metres. With the ACORK scientific team participating remotely from the Pacific Geoscience Centre in Sidney, BC, the ACORK data were downloaded and real-time data were collected for over an hour. This proved that the ACORK was functioning well and ready for next year's installation of additional equipment, including a tiltmeter. Following this test, a visual transect survey of the seafloor was completed and then the dive was wrapped up with the collection of a record 1-metre long push core by the ROPOS team!

ROPOS is currently heading to the seafloor to complete its second dive in the 889 area where equipment will be swapped out for repair, a new sonar will be installed to collect data in a dynamic area of gas release, and surveys of vents and gas expulsion at an area dubbed "bubbly gulch" will be collected. Tomorrow, after this dive and a CTD cast are completed, we travel to the 1027 area to begin the installation of the tsunami detection array.


June 2nd 1100 Update

02-Jun-12

The last of the Barkley instrument platforms was successfully installed on the NEPTUNE network this morning with the deployment of pod #4.  After adjusting the landing position of pod #4 and placing and connecting instruments on the seafloor from both pods, ROPOS returned to the ship to prepare for the next dive at a new location – ODP 889.  Underway by 1100, the R/V Thompson began steaming toward 889 with an anticipated arrival in the late afternoon.  The first dive at 889 will go to the advanced circulation obviation retrofit kit (ACORK) installed in Hole 1364A. ACORK 1364A, which extends 300 m below the seafloor, is equipped with five pressure gauges that monitor fluid pressures at the seafloor and below at sub-seabed depths of 155, 205, 245 and 295 metres. The dive objectives  include connecting the ACORK to NEPTUNE's network, testing the instruments, and downloading these pressure data.


June 2nd 0200 Update

02-Jun-12

In the final minutes of June 1, the NEPTUNE and ROPOS teams wrapped up a successful installation of the newly refurbished pod #1.  This pod was re-deployed in the axis of Barkley canyon with a full suite of instruments: ADCP, aquadop, sonar, and colour video camera.  After the placement of the equipment, ROPOS completed a video survey of the seafloor, mapping the benthic fauna along a depth transect from pod #1 to the Barkley hydrate instrument platform, located in shallower water.  This transect forms a baseline for future, repeat surveys to understand changes in time and with water depth.


June 1st 1600 Update

01-Jun-12

Early this morning, ROPOS returned to the ship after completing the deployment of pod #3 and the recovery of pod #4 in the mid-canyon area of Barkley. Upon recovery, the engineers found that there was a serious failure of the pan and tilt camera connector cable on pod #4 and the potential for it to be repaired onboard is limited. We have contacted the science team that works in this area to help us decide the best way forward for re-deploying pod #4. During the dive, repairs continued on pod #1, in preparation for its re-deployment.

After the pod dive today, a CTD cast, to measure water column properties and collect water samples, was completed at the mid-canyon area, and ROPOS was re-deployed to conduct a high resolution multibeam survey of the Barkley hydrate mound area. This survey will serve as a baseline for repeat surveys conducted in the future to assess the frequency, timing, size, and distribution of hydrate mound formation and collapse, which is ubiquitous there. The ROPOS multibeam survey is currently underway and targeted for completion this evening.

We plan next to return pod #1 to the axis of Barkley canyon with ROPOS, starting later this evening. After that ROPOS will be used to investigate Ogopogo, the vertical profiler system, to assess why its buoy won't release. A decision will be made, based on the status of the tsunami connector can builds and the pod #4 camera, to go to 889 or 1027 after the pod #1 dive.


May 31st 2300 Update

31-May-12

The Wally swap dive finished early this evening with good success. During the twelve hour ROPOS dive, the new Wally (Wally I) was deployed and the old Wally (Wally II) recovered. After a back and forth communication with NEPTUNE systems to rectify an initial ground fault, Laurenz Thomson in Bremen, Germany drove Wally I from its deployment site with lights and camera fully operational.

Two science experiments were maintained in Wally’s study area. For one experiment, investigating how carbonate persists among different marine environments, strings of mussel shells were deployed for long-term monitoring by Wally.

The second experiment was the recovery of a more than two-year long deployment of an osmotic water sampler. This instrument collects fluids over time by sucking seawater from the seafloor into a 300 m long coil of very thin tubing. Once the sampler is retrieved, the coils are unraveled, and discreet sections of the tubing are cut and analyzed to reconstruct a time series of fluid chemistry.

At midnight tonight, the deployment of pod #3 and recovery of pod #4 at Barkley will begin. The current plan is to complete all of the pod deployments at Barkley before transiting to the 1027 node area.


May 31st 1500 Update

31-May-12

The Wally swap dive started at~ 0630 this morning and is still underway. So far, Wally II, was disconnected and moved to ROPOS’ tool basket. As well, the osmotic sampler was recovered and stored in the basket. Wally I (the replacement for Wally II) was placed at a good landing sight at parking spot eighteen in Wallyland. Additional science was enabled by installing new mussel shell experiments and collecting cores.

Wally I’s cable is currently being uncoiled and “flown” from its spooled platform at parking spot eighteen to the Barkley hydrate instrument platform for connection to NEPTUNE’s power and internet. The new Wally will soon be driving around Wallyland.


May 30th 2300 Update

30-May-12

An ambitious three dives were planned and completed today by the extremely capable ROPOS team and the science and engineering teams. As reported earlier today, we recovered Pod #1 from the axis of Barkley channel for repair and re-deployment later in the cruise and deployed Pod #2 at the upper slope of Barkley canyon. Late this evening, we completed the recovery dive of Pod #3 (from the mid-east section of Barkley canyon) after conducting a video survey of the pod area, and collecting sediment core samples, a water sample, and a star fish that will help improve the confidence level of this species documented in NEPTUNE's new Marine Life guide. A busy day indeed.

Up next is a short multibeam survey while ROPOS and Wally I are prepared for the next dive. We anticipate that the Wally swap dive will be underway shortly after 0400 tomorrow, May 31st.


May 30th 1700 Update

30-May-12

Pod #2 was successfully deployed today at the upper slope site of Barkley canyon. Pod #2 comprises cameras, a sector scanning sonar, 2 ADCPs, and a sediment trap. During the deployment, camera and sector scanning sonar surveys of the site were completed. Sediment cores, water samples, and a sample of an anemone were recovered before ROPOS returned to the surface early this afternoon.

Multibeam surveys were completed after the Pod #2 deployment dive and ROPOS is about to be launched again to recover Pod #3 from Barkley canyon for repair. We anticipate that this dive will be completed around midnight tonight.

We plan to start May 31st with a ROPOS dive to swap out the Wallies at the Barkley hydrate area. The Wally dive should be completed in approximately 12 hours.


May 30th 0900 Update

30-May-12

Pod #1 was successfully recovered for repair from the axis of Barkley canyon. Prior to recovery, sediment push cores and a water sample were collected. Once on deck and ROPOS secured, multibeam surveys were continued while Pod #2 and ROPOS were readied for the next dive.

At ~2200 last night, ROPOS was ready to dive with Pod #2 latched, but during final dive checks, an error was identified in the fibre gyro-generated heading for ROPOS. Position information is critical for ROPOS to successfully conduct operations and a decision was made to delay the dive to repair ROPOS. Repairs continued until early this morning and the dive with Pod #2 began at ~0700 this morning. This dive is currently underway.

Next up today are: a short (2 hour) multibeam survey that will return the ship to the deeper locations in Barkley Canyon for a dive to recover Pod #3.


May 29 Afternoon Update

29-May-12

After completion of the second ROPOS dive this morning, multibeam surveys were run in an area northwest of the Barkley node where no high resolution bathymetry had yet been collected. After the survey, a CTD was run at the Barkley axis location, followed by the third ROPOS dive, currently underway. The current dive is at the canyon axis location where Pod 1 will be recovered for maintenance after seafloor video surveys and samples are collected.


May 29 AM Update

29-May-12

After completion of the successful first ROPOS dive yesterday morning, the afternoon operations included a CTD cast at the Barkley node location and a multibeam survey northwest of the Barkley area. The second ROPOS dive commenced late in the evening and continued this morning.

ROPOS operations began by deploying an instrument platform with a bottom pressure recorder onboard. The platform was then connected to the node via the recently repaired cable. NEPTUNE’s shorebased systems team found that the connection had no faults, which confirmed that the cable repair completed on the C/S Wave Venture earlier this month was successful.

Next, ROPOS connected Ogopogo, the vertical profiling system, to the instrument platform. Ogo powered up with no fault, but there are errors in the instructions for initiating profiling, which the systems team is currently de-bugging. The dive continued with connection to the already installed seismometer and deployment and connection of a bottom pressure recorder after adjusting the position of the instrument platform.

ROPOS completed the dive by conducting a patch test for multibeam calibration where the multibeam team measured vehicle pitch, role, and yaw along repeated passes of the same seafloor area.

Next up is a multibeam survey to capture new seafloor bathymetry and to gather water column data for bubble detection in the hydrate area where methane is currently being released at the Barkley hydrates area.


Ogopogo Deployed - May 28th 15:45 Update

28-May-12

Operations continue in the vicinity of the Barkley node. The first dive, which started this morning at ~0800, was on bottom at the node 1 hour later. The ROPOS team successfully connected the cable, recently repaired on the C/S Wave Venture, to the node and recovered the cable parking hardware to the surface.

Immediately following the recovery of ROPOS, the crew and NEPTUNE team members began preparations for the launch of the vertical profiling system, dubbed Ogopogo. This operation was methodically undertaken because of Ogo's large size and was successfully completed at ~1545 after it was acoustically released from the deployment cable and winch. ROPOS is currently being prepped for deploying the instrument platform that will connect Ogopogo, an already deployed seismometer, and a bottom pressure recorder to the NEPTUNE internet and power infrastructure.


Departure from Victoria - May 28th 12:25

27-May-12

The gangway was lifted and we set sail onboard the R/V Thomas G. Thompson toward Barkley Canyon at 1800 hours for the first cruise of the season – Wiring the Abyss 2012. The 37-person NEPTUNE team includes NEPTUNE staff, the Canadian Scientific Submersible Facility’s remotely operated vehicle, ROPOS, engineers from specialized companies, and a teacher. Our estimated time of arrival at Barkley is 0530 tomorrow morning where we will connect the newly repaired cable to the Barkley node, conduct tests of the multibeam system on ROPOS, and begin installation of the vertical profiling system, which we have dubbed Ogopogo because of its unique ability to move up and down in the full water column making measurements as it goes.

On our way to the first site, the NEPTUNE team was briefed by Captain John Wilson and other crew members on health, safety and the environment. After these important meetings were completed, the intense work continued preparing for the first ROPOS dive while everyone began to acclimate, in their own ways, their internal biological sleep clocks to adjust to life on 12 hours on/12 hours off and 4 hours on/8 hours off shifts.


Barkley Upper Slope readied - May 14th 08:35 Update

14-May-12

Over the past weekend the damaged extension cable that extends from Barkley Canyon node to the Upper Slope study area was re-terminated. This cable was capped and sealed last summer after the network sustained damage last year, likely from trawling activities.

The newly re-terminated cable will be connected to the Upper Slope instrument platforms by a jumper cable next month, and if power and communications are established, the remaining equipment will be deployed.


17 Days Until Cruise Begins - May 11 02:35 Update

11-May-12

Our June 2012 cruise begins in 17 days, and the people of NEPTUNE Canada are very busy preparing the equipment, plans and a new website to track cruise progress.


News home.